# Orak: A Foundational Benchmark for Training and Evaluating LLM Agents on Diverse Video Games

### v1

@@ 5.6 Effect of Fine-tuning
Model Finetune SF3 SuperMario DarkestD StarCraft2 SlaySpire BabaIsYou
Llama-3.2-1B 0.0$\pm$0.0 15.6$\pm$7.2 0.0$\pm$0.0 0.0$\pm$0.0 0.0$\pm$0.0 20.0$\pm$0.0
✓ 42.0$\pm$16.4 15.7$\pm$9.0 93.4$\pm$2.6 0.0$\pm$0.0 8.0$\pm$3.5 20.0$\pm$0.0
Llama-3.2-3B 12.0$\pm$11.0 13.7$\pm$4.3 87.2$\pm$9.5 0.0$\pm$0.0 0.0$\pm$0.0 20.0$\pm$0.0
✓ 40.0$\pm$7.1 17.2$\pm$5.7 92.0$\pm$0.2 0.0$\pm$0.0 10.7$\pm$1.2 20.0$\pm$0.0
GPT-4o 10.0$\pm$0.0 18.9$\pm$3.8 94.7$\pm$0.7 75.0$\pm$50.0 48.7$\pm$0.0 20.0$\pm$0.0
Table 7: Intra-game generalization accuracy.
Model Finetune SuperMario 2048
Llama-3.2-1B 18.7$\pm$8.6 0.1$\pm$0.1
✓ 26.7$\pm$12.3 2.8$\pm$1.8
Llama-3.2-3B 31.8$\pm$10.1 0.1$\pm$0.2
✓ 34.4$\pm$7.0 3.1$\pm$2.5
GPT-4o ✗ 34.1$\pm$14.2 7.0$\pm$5.7
Table 8: OOD-game accuracy.
Setup. We study the generalization capabilities of LLMs fine-tuned on expert trajectories described in Section 4. We consider two types of generalization: intra-game and OOD-game. Intra-game generalization evaluates whether an LLM can adapt to unseen scenarios within the same game, e.g., new stages, characters, or maps, that were not part of the training data. For this, we fine-tune the LLM on expert trajectories from all 12 games and evaluate it on 6 games that naturally provide unseen scenarios. Note that the scores are not directly comparable with those in Table 3, as the evaluation scenarios differ. In contrast, OOD-game generalization evaluates whether fine-tuning on a specific set of game trajectories enables the model to perform better on a held-out set of unseen games that were not encountered during training. Specifically, we hold out Super Mario and 2048, and fine-tune the model on trajectories from the remaining 10 games. We used the same evaluation scenario as Table 3. See Appendix P for details on the unseen scenarios used in intra-game generalization experiments.
@@
Intra-game Generalization. Table 8 shows performance of fine-tuned Llama-3.2-1B/3B models on unseen scenarios, compared to the pretrained ones. Despite their relatively smaller scale, we find that these models demonstrate strong generalization to unseen scenarios; in 4 out of 6 games, both fine-tuned models outperform their pretrained counterparts. The performance gain largely comes from the model learning to generate valid actions more reliably after fine-tuning, especially in environments where the pretrained model frequently failed to act meaningfully. However, this approach appears insufficient to significantly enhance the spatial reasoning abilities of smaller models. This is evident in Baba Is You, where fine-tuned models struggle to construct the winning condition required to score above 20.0—a task that remains challenging even for GPT-4o.
@@
OOD-game Generalization. Table 8 shows OOD-game generalization performance of Llama-3.2-1B/3B models. Despite being trained exclusively on trajectories from different games, the performance of these models on Super Mario and 2048 improves significantly. This suggests that fine-tuning on trajectories shaped by reflection and planning enables models to learn transferable decision-making routines, as the underlying capabilities required for these behaviors are shared across games, even when the games themselves differ. Specifically, the improvement in 2048 likely benefits from its structural similarity to Baba Is You, a training game that also uses a 2D grid layout and discrete directional actions (up, down, left, right), which may have facilitated generalization.
6 Conclusion
@@
In this paper, we introduce Orak, a benchmark designed to train and evaluate LLM agents across diverse real-world video games. Our benchmark enables comprehensive assessments of LLM capabilities required to play most game genres. Through a plug-and-play interface powered by MCP, it further allows consistent evaluation of rapidly evolving LLMs over various agentic modules. In addition, we release a fine-tuning dataset, consisting of game interaction trajectories of top-performing LLMs, which can effectively transform pre-trained LLMs into gaming agents. With the comprehensive game set and user-friendly interface, Orak sets a new foundation for game-based LLM evaluation, driving progress towards versatile and high-performing gaming agents.
@@ Appendix A Game Genre Categorization
Video games are generally categorized into six widely recognized genres, i.e., Action, Adventure, Role-Playing, Simulation, Strategy, and Puzzle, each characterized by distinct gameplay structures [73]. Action games emphasize responsiveness and precise physical control. Adventure games focus on narrative exploration, interactive dialogues, and clue-based progression grounded in logical inference. Role-Playing games focus on character progression, stat-driven combat mechanics, and quest-based narrative development. Simulation games present system-driven environments in which players manage complex and interdependent variables such as time, resources, and procedural systems. Strategy games are a genre that emphasizes planning, resource management, decision-making, and tactical execution. Puzzle games revolve around rule-based problem solving, pattern recognition, and logical or spatial reasoning, typically within a clearly defined system of constraints.
@@
(a) Street Fighter III [53] is classified as an Action game, as it primarily relies on responsiveness, frame-precise input, and physical dexterity. Its gameplay requires fast reflexes and mastery of complex input sequences.
@@
(b) Super Mario [54] is categorized as Action game. The core gameplay emphasizes precise timing in jumping and movement, demanding moment-to-moment control in response to environmental hazards and enemy placements.
@@
(c) Ace Attorney [55] is labeled as an Adventure game due to its narrative-driven structure, reliance on clue collection, and logical deduction. Players progress by interacting with characters and uncovering story elements through investigative mechanics, with minimal emphasis on reflex-based input.
@@
(d) Her Story [56] similarly fits within the Adventure category. Though more experimental in form, it shares a strong focus on narrative discovery through a search-based interface, requiring players to piece together a fragmented story using non-linear exploration and deductive reasoning.
@@
(e) Pokémon Red [57] is classified as a Role-Playing game, as its primary mechanics involve turn-based combat, character progression, stat management, and quest-driven exploration. Additionally, Adventure was assigned as a secondary genre to reflect the game’s emphasis on world exploration, interaction with non-player characters, and sequential progression through narrative landmarks.
@@
(f) Darkest Dungeon [58] is similarly assigned Role-Playing as the main genre, supported by stat-driven character development and progression systems. However, it also includes significant Strategy elements, as players must carefully manage resources, form party compositions, and make tactical decisions in turn-based combat. Adventure was further added as a secondary genre due to its dungeon-crawling structure and emphasis on risk-driven exploration.
@@
(g) Minecraft [59] is categorized as a Simulation game due to its open-ended, system-driven mechanics, including resource gathering, crafting, and environmental manipulation. It also exhibits Role-Playing traits through its progression systems and player-driven narrative development, warranting secondary classification.
@@
(h) Stardew Valley [61] is also assigned to the Simulation genre based on its emphasis on time management, farming systems, and interrelated mechanics spanning multiple in-game variables. It incorporates Role-Playing through relationship-building and character progression, and Adventure through dungeon exploration, seasonal events, and quest-based interactions.
@@
(i) StarCraft II [62] is classified as a Strategy game, consistent with its real-time strategic planning, resource allocation, and micromanagement mechanics. Given the importance of unit-level control, Action was added as a secondary genre to reflect the real-time, reflex-driven demands during gameplay.
@@
(j) Slay the Spire [63] is categorized solely as a Strategy game. The gameplay centers around deck-building, route optimization, and turn-based combat, requiring players to plan several moves.
@@
(k) Baba Is You [64] is clearly identified as a Puzzle game, as its mechanics are centered on solving logic-based problems through the manipulation of in-game rules represented by words. The core loop involves constrained, rule-based problem solving and spatial reasoning.
@@
(l) 2048 [65] is also classified as a Puzzle game, characterized by deterministic mechanics, arithmetic pattern recognition, and constraint-based spatial logic within a fixed grid.
@@ C.1 Game Description for Street Fighter III
Environment. Street Fighter III [53] is a 2D competitive fighting game, known for precise controls, deep mechanics, and a diverse roster of characters. Each character features unique moves, combos, and super arts, requiring precise timing and strategic decision-making. Players aim to defeat their opponent through a mix of normal attacks, special moves, and advanced mechanics like parries and cancels. For implementation, we use Diambra Arena environment [74], a Docker-based platform designed for RL research. Street Fighter III is one of the environments supported by Diambra, which not only enables seamless extraction of the game state—such as screenshots, health, timer, and super bar values—but also provides a straightforward interface for sending controller inputs. In this setting, the agent operates in a discrete action space that directly corresponds to various controller actions, including directional movement, attack buttons, and their combinations, enabling intuitive and fine-grained control of the in-game character.
@@
Figure 6: Character detection using YOLOv11 model [75] in Street Fighter III.
Observation-to-Text Conversion. The Diambra environment offers a convenient interface for extracting the game state from Street Fighter III. Through this interface, we obtain the latest game frame at a resolution of 224×384, along with key state information such as remaining time, player and opponent health, super bar gauge, super count, stun bar gauge, and stun status. However, a critical aspect in fighting games is understanding the relative positions of the characters, which is not directly provided by the Diambra environment. To address this limitation, we employ a lightweight YOLOv11 object detection model [75] to extract the relative positions of the two characters from the game frame. Based on the computed distance, we classify the spatial relationship into three discrete categories—very close, close, and far—and incorporate this information into the agent’s user prompt.
@@
  • Movement actions (8 total): Left, Left$+$Up, Up, Up$+$Right, Right, Right$+$Down, Down, Down$+$Left
  • Attack actions (9 total): Low Punch, Medium Punch, High Punch, Low Kick, Medium Kick, High Kick, Low Punch$+$Low Kick, Medium Punch$+$Medium Kick, High Punch$+$High Kick
@@
  • Character-agnostic actions (14 total): Move Closer, Move Away, Jump Closer, Jump Away, Super Attack, Low Punch, Medium Punch, High Punch, Low Kick, Medium Kick, High Kick, Low Punch$+$Low Kick, Medium Punch$+$Medium Kick, High Punch$+$High Kick
@@
For example, if the character is positioned on the left side of the screen and the high-level action is ‘Move Closer’, the system issues ‘Right’ movement commands over four frames. If the action is ‘Fireball’, the corresponding low-level sequence would be ‘Down’$\rightarrow$‘Down$+$Right’$\rightarrow$‘Right’$\rightarrow$‘Medium Punch’. Since the number of character-specific actions varies, the total size of the high-level action space differs depending on character, typically ranging around 20 actions.
@@ C.3 Evaluation Metric for Street Fighter III
[…]
@@
We adopt a Bradley-Terry model formulation [76] for Elo estimation, where each model’s rating is iteratively optimized using gradient ascent on the log-likelihood of observed outcomes. The gradient is computed based on the expected win probabilities derived from current ratings, using the standard Elo transformation:
@@ C.5 Result for Street Fighter III
Random Agent 10.0$\pm$6.4 12
Llama-3.2-1B 0.0$\pm$0.0 13
Llama-3.2-3B 13.3$\pm$5.8 10.5
Qwen-2.5-3B 20.0$\pm$0.0 4.5
Qwen-2.5-7B 16.7$\pm$11.5 7.5
Minitron-4B 16.7$\pm$11.5 7.5
Minitron-8B 23.3$\pm$5.8 3
GPT-4o-mini 16.7$\pm$11.5 7.5
GPT-4o 29.7$\pm$14.3 2
o3-mini 33.3$\pm$15.3 1
Gemini-2.5-pro 13.3$\pm$11.5 10.5
Claude-3.7 16.7$\pm$11.5 7.5
Deepseek-R1 20.0$\pm$0.0 4.5
@@
Random Agent - 10.0$\pm$6.4 9
Llama-3B Zero-shot 13.3$\pm$5.8 8
Reflection 30.0$\pm$17.3 1.5
Planning 20.0$\pm$0.0 6
Ref-Plan 16.7$\pm$20.8 7
GPT-4o Zero-shot 29.7$\pm$14.3 3
Reflection 23.3$\pm$20.8 4.5
Planning 30.0$\pm$26.5 1.5
Ref-Plan 23.3$\pm$20.8 4.5
@@
Random Agent - 10.0$\pm$6.4 10
GPT-4o Text 29.7$\pm$14.3 2
Image 23.7$\pm$15.9 4
Both 24.3$\pm$14.5 3
Gemini Text 13.3$\pm$11.5 9
Image 16.7$\pm$11.5 7
Both 20.0$\pm$10.0 6
Claude Text 16.7$\pm$11.5 7
Image 23.3$\pm$11.5 5
Both 33.3$\pm$5.8 1
@@ D.1 Game Description for Super Mario
Environment. Super Mario (1985 Super Mario Bros) [54] is a side-scrolling game where the player controls Mario to avoid obstacles, defeat monsters, and reach the flag. In this environment, Mario progresses through the game using directional key controls (e.g., ‘left’ and ‘right’ keys) and jump actions. Mario should either destroy or traverse obstacles (e.g., bricks, stairs, pipes), avoid or defeat monsters (e.g., Goombas, Koopas) by jumping on them and avoid falling into pits. For implementation, we adopt the gym_super_mario_bros environment [77], which is widely used in the reinforcement learning (RL) community. Specifically, we use the SuperMarioBros-v1 environment, where Mario plays within a 256$\times$240-pixel screen with a black background, as shown in Figure 10(a). The environment consists of 8 worlds, each containing 4 stages. We use World 1, Stage 1 as our default evaluation setting (for Table 3). However, the environment supports evaluation across any world and stage. For example, we use World 3, Stage 1 for the evaluation of intra-game generalization (for Table 8).

### v2

@@ 5.6 Effect of Fine-tuning
Models Finetune Intra-Game OOD-Game Non-Game
SF3 DarkestD StarCraft2 SlaySpire BabaIsYou SuperMario 2048 Math500 WebShop-E WebShop-H
[…]
Table 7: Generalization performance of LLMs fine-tuned on expert gameplay trajectories from Orak.
Setup.
To study the effect of fine-tuning, we consider three types of generalization: Intra-game, OOD-game, and Non-game. Intra-game generalization evaluates whether an LLM can adapt to unseen scenarios within the same game, e.g., new stages or characters. OOD-game generalization evaluates whether fine-tuning on a specific set of games enables the model to act better on unseen games. Non-game generalization evaluates whether our fine-tuning set with agent gameplay trajectory helps the model to perform better on non-game tasks like math or web navigation. For Intra-game and Non-game generalizations, we fine-tune LLMs on all 12 games, while for OOD-game generalization, we split 12 games in to 10 training games and 2 test games, i.e., Super Mario and 2048. See Appendix P for detailed fine-tuning configurations and unseen scenario details in intra-game generalization.
Intra-game Generalization.
@@
As shown in Table 7, fine-tuned Llama-3.2-1B/3B show better generalization to unseen scenarios than the pretrained ones; in 3 out of 5 games, both fine-tuned models outperform their pretrained counterparts. The performance gain largely comes from the model learning to generate valid actions more reliably after fine-tuning, especially in environments where the pretrained model frequently failed to act meaningfully. However, this approach appears insufficient to significantly enhance the spatial reasoning abilities of smaller models. This is evident in Baba Is You, where fine-tuned models struggle to construct the winning condition required to score above 20.0.
OOD-game Generalization.
@@
Similarly, despite being trained exclusively on trajectories from different games, fine-tuned models perform significantly better on OOD games like Super Mario and 2048. This suggests that fine-tuning on trajectories shaped by reflection and planning enables models to learn transferable decision-making routines, as the underlying capabilities required for these behaviors are shared across games. Specifically, the improvement in 2048 likely benefits from its structural similarity to Baba Is You, a training game that also uses 2D grid layout and discrete move actions (up, down, left, right). More fine-tuning analysis with varying OOD gaps are in Appendix Q.
Non-game Generalization.
[…]
6 Discussion
Conclusion.
@@
In this paper, we introduce Orak, a benchmark designed to train and evaluate LLM agents across diverse video games. Orak enables comprehensive assessments of LLM capabilities required to play most game genres. Through a plug-and-play interface powered by MCP, it allows consistent evaluation of rapidly evolving LLMs over various agentic modules. In addition, we release a fine-tuning dataset of game interaction trajectories of top-performing LLMs, which can effectively transform pre-trained LLMs into gaming agents. With the comprehensive game set and user-friendly interface, Orak sets a new foundation for game-based LLM evaluation, driving progress towards versatile and high-performing gaming agents.
Limitations.
We note that Orak provides environments favorable for LLM reasoning and decision-making by pre-processing game states into structured text where information unnecessary to gameplay is hidden. This may offer insights for providers seeking to deploy resource-efficient LLMs on games. We leave a complementary direction, providing the full in-game states with rich uncurated texts to LLMs, for future work. Further discussions, including cost and licensing, are provided in Appendix R.
@@ Appendix A Game Genre Categorization
Video games are generally categorized into six widely recognized genres, i.e., Action, Adventure, Role-Playing, Simulation, Strategy, and Puzzle, each characterized by distinct gameplay structures (Wikipedia contributors, 2025). Action games emphasize responsiveness and precise physical control. Adventure games focus on narrative exploration, interactive dialogues, and clue-based progression grounded in logical inference. Role-Playing games focus on character progression, stat-driven combat mechanics, and quest-based narrative development. Simulation games present system-driven environments in which players manage complex and interdependent variables such as time, resources, and procedural systems. Strategy games are a genre that emphasizes planning, resource management, decision-making, and tactical execution. Puzzle games revolve around rule-based problem solving, pattern recognition, and logical or spatial reasoning, typically within a clearly defined system of constraints.
@@
(a) Street Fighter III (Capcom, 1997) is classified as an Action game, as it primarily relies on responsiveness, frame-precise input, and physical dexterity. Its gameplay requires fast reflexes and mastery of complex input sequences.
@@
(b) Super Mario (Christian Kauten, 2018) is categorized as Action game. The core gameplay emphasizes precise timing in jumping and movement, demanding moment-to-moment control in response to environmental hazards and enemy placements.
@@
(c) Ace Attorney (Capcom, 2001) is labeled as an Adventure game due to its narrative-driven structure, reliance on clue collection, and logical deduction. Players progress by interacting with characters and uncovering story elements through investigative mechanics, with minimal emphasis on reflex-based input.
@@
(d) Her Story (Barlow, 2015) similarly fits within the Adventure category. Though more experimental in form, it shares a strong focus on narrative discovery through a search-based interface, requiring players to piece together a fragmented story using non-linear exploration and deductive reasoning.
@@
(e) Pokémon Red (Game Freak, 1996) is classified as a Role-Playing game, as its primary mechanics involve turn-based combat, character progression, stat management, and quest-driven exploration. Additionally, Adventure was assigned as a secondary genre to reflect the game’s emphasis on world exploration, interaction with non-player characters, and sequential progression through narrative landmarks.
@@
(f) Darkest Dungeon (Red Hook Studios, 2016) is similarly assigned Role-Playing as the main genre, supported by stat-driven character development and progression systems. However, it also includes significant Strategy elements, as players must carefully manage resources, form party compositions, and make tactical decisions in turn-based combat. Adventure was further added as a secondary genre due to its dungeon-crawling structure and emphasis on risk-driven exploration.
@@
(g) Minecraft (Mojang Studios, 2011) is categorized as a Simulation game due to its open-ended, system-driven mechanics, including resource gathering, crafting, and environmental manipulation. It also exhibits Role-Playing traits through its progression systems and player-driven narrative development, warranting secondary classification.
@@
(h) Stardew Valley (ConcernedApe, 2016) is also assigned to the Simulation genre based on its emphasis on time management, farming systems, and interrelated mechanics spanning multiple in-game variables. It incorporates Role-Playing through relationship-building and character progression, and Adventure through dungeon exploration, seasonal events, and quest-based interactions.
@@
(i) StarCraft II (Blizzard Entertainment, 2010) is classified as a Strategy game, consistent with its real-time strategic planning, resource allocation, and micromanagement mechanics. Given the importance of unit-level control, Action was added as a secondary genre to reflect the real-time, reflex-driven demands during gameplay.
@@
(j) Slay the Spire (MegaCrit, 2017) is categorized solely as a Strategy game. The gameplay centers around deck-building, route optimization, and turn-based combat, requiring players to plan several moves.
@@
(k) Baba Is You (Hempuli, 2019) is clearly identified as a Puzzle game, as its mechanics are centered on solving logic-based problems through the manipulation of in-game rules represented by words. The core loop involves constrained, rule-based problem solving and spatial reasoning.
@@
(l) 2048 (Cirulli, 2014) is also classified as a Puzzle game, characterized by deterministic mechanics, arithmetic pattern recognition, and constraint-based spatial logic within a fixed grid.
@@ C.1 Game Description for Street Fighter III
Environment. Street Fighter III (Capcom, 1997) is a 2D competitive fighting game, known for precise controls, deep mechanics, and a diverse roster of characters. Each character features unique moves, combos, and super arts, requiring precise timing and strategic decision-making. Players aim to defeat their opponent through a mix of normal attacks, special moves, and advanced mechanics like parries and cancels. For implementation, we use Diambra Arena environment (DIAMBRA, 2025), a Docker-based platform designed for RL research. Street Fighter III is one of the environments supported by Diambra, which not only enables seamless extraction of the game state—such as screenshots, health, timer, and super bar values—but also provides a straightforward interface for sending controller inputs. In this setting, the agent operates in a discrete action space that directly corresponds to various controller actions, including directional movement, attack buttons, and their combinations, enabling intuitive and fine-grained control of the in-game character.
@@
Figure 6: Character detection using YOLOv11 model (Khanam & Hussain, 2024) in Street Fighter III.
Observation-to-Text Conversion. The Diambra environment offers a convenient interface for extracting the game state from Street Fighter III. Through this interface, we obtain the latest game frame at a resolution of 224×384, along with key state information such as remaining time, player and opponent health, super bar gauge, super count, stun bar gauge, and stun status. However, a critical aspect in fighting games is understanding the relative positions of the characters, which is not directly provided by the Diambra environment. To address this limitation, we employ a lightweight YOLOv11 object detection model (Khanam & Hussain, 2024) to extract the relative positions of the two characters from the game frame. Based on the computed distance, we classify the spatial relationship into three discrete categories—very close, close, and far—and incorporate this information into the agent’s user prompt.
@@
  • Movement actions (8 total): Left, Left++Up, Up, Up++Right, Right, Right++Down, Down, Down++Left
  • Attack actions (9 total): Low Punch, Medium Punch, High Punch, Low Kick, Medium Kick, High Kick, Low Punch++Low Kick, Medium Punch++Medium Kick, High Punch++High Kick
@@
  • Character-agnostic actions (14 total): Move Closer, Move Away, Jump Closer, Jump Away, Super Attack, Low Punch, Medium Punch, High Punch, Low Kick, Medium Kick, High Kick, Low Punch++Low Kick, Medium Punch++Medium Kick, High Punch++High Kick
@@
For example, if the character is positioned on the left side of the screen and the high-level action is ‘Move Closer’, the system issues ‘Right’ movement commands over four frames. If the action is ‘Fireball’, the corresponding low-level sequence would be $\rightarrow‘Down$$\rightarrow‘Right$$\rightarrow$‘Medium Punch’. Since the number of character-specific actions varies, the total size of the high-level action space differs depending on character, typically ranging around 20 actions.
@@ C.3 Evaluation Metric for Street Fighter III
$\text{Score}=\text{Number of stages cleared by the agent}\times 10.$
@@
We adopt a Bradley-Terry model formulation (Bradley & Terry, 1952) for Elo estimation, where each model’s rating is iteratively optimized using gradient ascent on the log-likelihood of observed outcomes. The gradient is computed based on the expected win probabilities derived from current ratings, using the standard Elo transformation:
@@ C.5 Result for Street Fighter III
[…]
Llama-3.2-1B $\pm0.0$ 13
[…]
Gemini-2.5-pro $3±\pm1$1.5 10.5
[…]
@@
[…]
Reflection $3±\pm2$0.8 4.5
[…]
Ref-Plan $3±\pm2$0.8 4.5
@@
[…]
Gemini Text $3±\pm1$1.5 9
[…]
@@ D.1 Game Description for Super Mario
[…]